For a projectile moving along a parabolic path, the maximum height attained may be:
1
Expert's answer
2020-11-20T09:29:53-0500

The maximum height for a body launched at an angle to the horizontal, is:


"h_{max} = \dfrac{v^2\sin^2\theta}{2g}"

where "v" is the initial speed of the body, "\theta" is launching angle and "g = 9.81m\/s^2" is the gravitational acceleration.

See https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/ for more details.


Answer. "h_{max} = \dfrac{v^2\sin^2\theta}{2g}".